OpenStudy (anonymous):

Match the information about interest earned with the principal amount invested. Remember the equation for calculating simple interest is I equals p times r times t. Q)forty−two dollars over 2 years at an interest rate of three percent A) p equals five hundred dollars B) p equals seven hundred dollars C) p equals four hundred dollars D) p equals six hundred dollars

OpenStudy (nurali):

I = PRT Where I = 42 Dollar, R = 3% = 0.03 and T = 2 year 42 = P * 0.03 * 2 42 = 0.06P 42/0.06 = P P = 700$